An ice cube slides from rest without friction down a long inclined ramp that makes an angle of 39.0 o with the horizontal. What is the speed of the cube after it slides 1.97 m down the plane?
Answer #1

V^2 = U^2 + 2aS

a = gsin(theta)

=> V^2 = 0 + 2*9.8*sin(39)*1.97

=> V = sqrt (2*9.8*0.629*1.97) = 4.92 m/s after it slides 1.97 m down the plane